# Susan Athey    Emil Palikot

### v4

@@ 5.1 Average Treatment Effects
The right panel decomposes selection from treatment effects in Mentoring. Applicants selected by mentors but randomized to control achieve substantially higher technology employment (43% at month thirteen) than non-selected applicants (19%), confirming that mentors identify promising candidates. Furthermore, the treatment-control gap among selected applicants is smaller than the selection effect itself, underscoring the importance of randomization for isolating program impacts from mentor selection.
@@
Note: Share of subjects with a tech job per group across time. The month zero is the month of application to Dare IT. In the left figure, we group subjects according to the experimental groups: the treatment group in Mentoring (in green), the treatment group in Challenges (in red), and Out of Dare IT (in blue), where the last group consists of the applicants in control groups, non-selected applicants to Mentoring, and late applicants to Challenges. In the right figure, we group applicants into those selected and those not selected by mentors. "Selected and treated" group (in green) shows the outcomes of the treatment group in Mentoring, "Selected but not treated" group (in blue) is the control group of Mentoring "Not selected" group (in red) includes applicants to Mentoring that were not selected by mentors, both experimental groups of Challenges, and late applicants to Challenges.
Table 2 presents estimated average treatment effects. We estimate the averate treatment effects using the difference-in-means estimator (columns 1-3, 5-7) and using Cox proportional hazard models that exploit monthly outcome variation (columns 4, 8). Appendix F verifies the proportional hazards assumption; Appendix C presents Kaplan-Meier survival curves confirming the timing patterns in Figure 2.
Both programs generate substantial, statistically significant increases in technology employment. Mentoring increases the probability of technology employment by 15.3 percentage points (SE 5.7), representing a 34% increase relative to the control group baseline of 45.6%. Challenges produces an 11.3 percentage point effect (SE 4.8), a 42% increase relative to the 26.6% control baseline. Both estimates use difference-in-means; results are nearly identical using Cox proportional hazard models that exploit monthly variation in job start dates (14.5pp for Mentoring, 11.2pp for Challenges).
Table 2: Average Treatment Effects of Mentoring and Challenges.
Difference-in-means Prop. hazard Difference-in-means Prop. hazard
@@
New job Tech job Non-tech job Tech job Cox New job Tech job Non-tech job Tech job Cox
@@
SE (0.057) (0.057) (0.031) (0.055) (0.051) (0.048) (0.042) (0.051)
ATE % baseline 13.576 33.676** -65.641** 31.718*** 14.458 42.646** -17.600 42.232**
SE (9.893) (14.783) (26.720) (11.992) (10.268) (18.131) (17.978) (19.301)
No. obs. treated 151 151 151 2265 166 166 166 2656
No. obs. control 147 147 147 2205 218 218 218 3488
Note: The estimates of the average treatment effects of Mentoring and Challenges on the probability of finding a new job, a new job in technology, and a new job not in technology. Columns 1-3 and 5-7 present estimates using the difference-in-means estimator and columns 4 and 8 using a proportional hazard model with a full set of covariates. The first row shows estimates of the average treatment effect in the difference in the probability of reporting employment at month 15. The third row presents the estimates from the first row divided by the mean outcome in the control group. Standard errors are presented below the average treatment effect estimates. The variance of the difference-in-means estimate of the Challenges intervention is estimated using the classical Neyman-Splawa estimator. *** p<0.01, ** p<0.05, * p<0.1.
@@
Estimates in Column 3 of Table 2 suggest that Mentoring operates partly through increased job search selectivity. While Mentoring increases overall employment by only 7.8 percentage points (SE 5.7), statistically insignificant, it increases technology employment by 15.3 percentage points and decreases non-technology employment by 7.6 percentage points (SE 3.1). This implies that approximately half of Mentoring’s impact reflects participants becoming more selective, holding out for technology positions rather than accepting readily available employment in other sectors. In contrast, Challenges shows no significant effect on non-technology employment (-4.1pp, SE 4.2), consistent with this program operating through skill signaling rather than changing search behavior or persistence. The differential patterns support distinct mechanisms: Mentoring increases both participants’ capability to access technology jobs (through networks and tacit knowledge) and their willingness to persist in pursuing them (for example, through confidence-building), while Challenges primarily improves application success conditional on applying. This finding provides evidence consistent with the self-efficacy channel hypothesized in Section 3: mentors help women maintain focus on technology careers.
@@
In Appendix E, we show the results based on the outcomes survey. We find that the survey respondents in both treatment groups report a higher number of job offers and higher salaries than the control group. The differences between treatment and control are greater in the case of Mentoring than Challenges.
5.2 Impact of Program Participation on the Number of LinkedIn Connections
@@
We directly test the network mechanism by examining whether program participation increases professional network size, measured by LinkedIn connection counts. This analysis distinguishes network expansion from other channels: if Mentoring works solely through skill development or signaling, we should observe no differential effect on connections relative to Challenges. Conversely, if networking is a primary mechanism, Mentoring should increase connections while Challenges, which operates entirely online without mentor introductions or employer contact, should not.
@@
The data support the network mechanism. Mentoring increases LinkedIn connections by 52.7 (SE 20.8) from a control group baseline of 225, a 23% increase. In contrast, Challenges produces a statistically insignificant reduction of 11.4 connections (SE 17.8) from a baseline of 232. The differential treatment effect of 64 connections between programs (SE 28) provides direct evidence that Mentoring expands professional networks while Challenges does not.
@@
To understand where these effects appear in the connection distribution, Table 3 presents quantile regression estimates. Mentoring generates large, statistically significant increases at the 25th percentile (62 connections, SE 27) and median (92 connections, SE 35), but no detectable effect at the 75th percentile. This pattern indicates the program particularly benefits participants with initially sparse networks. In contrast, Challenges shows no significant effects at any quantile, confirming this program does not operate through networking channels.
Table 3: Treatment Effects on LinkedIn Professional Networks.
Quantile Treatment Effect Mentoring Treatment Effect Challenges Difference: Mentoring and Challenges
0.25 62** 5 57*
@@
Note: The table provides the quantile regression results for the Challenges and Mentoring programs. The treatment effects of the experimental groups on LinkedIn connections are shown at the 25th, 50th, and 75th quantiles. The difference column captures the difference in the treatment effects between the Mentoring and Challenges programs. Estimates are obtained by regressing the treatment indicator on the number of connections. Standard errors are in parentheses. *** p<0.01, ** p<0.05, * p<0.1.
The magnitude of network expansion is economically meaningful. A 92-connection increase at the median represents substantial growth from the control baseline of 225, comparable to the network size differences between mid-career and senior professionals in our sample. Moreover, the increase occurs despite LinkedIn’s 500-connection display cap, which likely attenuates our estimates: 25% of Mentoring treatment participants hit this ceiling compared to 23% of controls, suggesting the true effect exceeds our measured 52.7-connection average.
This network expansion cannot be solely attributed to new employment. Both Mentoring and Challenges participants secure technology jobs at high rates, yet only Mentoring participants expand their networks. Had job acquisition been the primary driver of connection growth, through adding new colleagues and industry contacts post-employment, we should observe comparable increases in both programs. The absence of connection growth among Challenges participants, despite their 38% technology employment rate, indicates that Mentoring’s network effects operate independently of (and plausibly prior to) job acquisition. This timing is consistent with mentors actively facilitating introductions during the program, as reported in our qualitative interviews, rather than networks expanding passively after job placement.
@@ 5.3 Heterogeneous Treatment Effects
The hypothesized contrasting mechanisms underlying Mentoring and Challenges generate testable predictions about which participants benefit most from each program. We exploit treatment effect heterogeneity to provide evidence consistent with these mechanisms: if Mentoring operates primarily through network expansion (as suggested by LinkedIn connections analysis), effects should concentrate among participants with weak existing networks; if Challenges works through skill signaling, effects should be largest for those with weaker pre-treatment credentials. We estimate heterogeneous treatment effects using Cox proportional hazard models that estimate treatment effects in subgroups defined by the pre-treatment characteristics of interest. We control for the full covariate set.
Table 4 presents heterogeneous treatment effects across three participant characteristics that map to credential strength and network access. These patterns provide evidence on which populations each program serves most effectively and tests the distinct mechanisms we hypothesize.
@@
The clearest differential pattern emerges from professional experience. Challenges generates 28 percentage points larger treatment effects for participants with under five years of experience compared to those with longer careers (SE 10), whereas Mentoring shows the opposite pattern with 20 percentage points larger effects for experienced participants (SE 11). The cross-program differential is 49 percentage points (SE 11). This contrast directly supports our theoretical predictions: early-career women lack established work history to credibly signal their technical abilities, making the Challenges portfolio particularly valuable as a complement to sparse employment records.
Educational background reveals similar patterns. Recall that all participants were pre-screened for technical competency, so heterogeneity by educational background likely reflects differences in credential signals and professional networks, and to a lesser degree underlying ability. Among Mentoring participants, those without STEM degrees show 19 percentage points larger effects than those with STEM backgrounds (SE 11). This pattern is consistent with the network access: non-STEM participants followed educational paths outside technology-adjacent fields, limiting opportunities to build connections with peers entering the sector.
In contrast, Challenges shows no STEM heterogeneity (difference: 0.3pp, SE 11). This null result demonstrates that portfolio signals are equally valuable regardless of educational credentials—employers evaluate demonstrated capability rather than inferring skill from degrees.
@@
Geography provides additional evidence on network mechanisms. Mentoring generates 26 percentage point treatment effects for participants from small towns (SE 10) compared to only 3 percentage points for Warsaw residents (difference: 23pp, SE 14, p=0.11). This pattern aligns with network barriers binding more tightly outside major urban centers, where local technology ecosystems are sparse and professional connections to the sector are limited. Mentor networks based in Warsaw and other major cities provide greater marginal value when participants lack local access to technology professionals. Challenges shows no significant geographic heterogeneity (difference: -2pp, SE 1.2), consistent with online portfolio signals functioning equally regardless of location, employers evaluate portfolios based on technical content rather than applicant geography.
Table 4: Heterogeneity in Treatment Effects: Mentoring vs. Challenges.
@@
Baseline CATE Baseline CATE Diff. Mentoring – Challenges
@@
Yes 0.538 (0.057) 0.038 (0.079) 0.283 (0.059) 0.12 (0.099) -0.081 (0.082)
No 0.362 (0.058) 0.228 (0.07) 0.259 (0.035) 0.116 (0.057) 0.111 (0.068)
Difference subgroups 0.176 (0.081) -0.189 (0.106) 0.024 (0.068) 0.003 (0.114) -0.193 (0.106)
Length of professional experience
More than 5 years 0.412 (0.049) 0.202 (0.066) 0.31 (0.039) -0.003 (0.065) 0.205 (0.063)
Less than 5 years 0.556 (0.075) 0 (0.087) 0.19 (0.042) 0.281 (0.075) -0.281 (0.086)
Difference subgroups -0.144 (0.089) 0.203 (0.109) 0.120 (0.057) -0.284 (0.099) 0.486 (0.106)
Small Town
Small town 0.429 (0.085) 0.26 (0.101) 0.256 (0.067) 0.149 (0.13) 0.111 (0.108)
Warsaw 0.447 (0.082) 0.033 (0.103) 0.182 (0.044) 0.166 (0.089) -0.133 (0.093)
Difference subgroups -0.019 (0.118) 0.227 (0.144) 0.074 (0.157) -0.019 (0.118) 0.246 (0.197)
Note: Estimates of the average treatment effects conditioned on subgroups in Mentoring and Challenges. The outcome variable is Tech job, and estimates are obtained using Cox proportional hazards models per subgroup with a full set of covariates. The first and third columns show the average outcome per subgroup within the control groups. The second and fourth columns show the difference between the treatment and control groups in each subgroup. The last column shows the difference in the treatment effects between Mentoring and Challenges. Standard errors are in parentheses.
Additional Characteristics from Mentoring.
The Mentoring registration survey was more extensive than the Challenges one, allowing us to observe more characteristics. We also obtain the characteristics of mentors from their registration survey. In Table 5, we present additional results documenting heterogeneity in treatment effects along these dimensions.
Table 5: Heterogeneity in Treatment Effects: Mentoring.
Characteristic variable value is 0 Characteristic variable value is 1
Baseline CATE Baseline CATE Diff. Subgroups
@@
Mothers 0.237 (0.07) 0.348 (0.097) 0.532 (0.048) 0.071 (0.062) 0.277 (0.115)
Applied before 0.5 (0.087) 0.139 (0.109) 0.442 (0.047) 0.143 (0.06) -0.004 (0.124)
@@
Managerial exp. 0.378 (0.056) 0.259 (0.073) 0.549 (0.059) 0.002 (0.074) 0.257 (0.104)
Long exp. in tech 0.39 (0.063) 0.277 (0.085) 0.512 (0.054) 0.021 (0.066) 0.257 (0.107)
1st time mentor 0.444 (0.062) 0.155 (0.08) 0.476 (0.055) 0.1 (0.073) 0.056 (0.108)
Note: Estimates of the average treatment effects conditioned on subgroups in the Mentoring program. The outcome variable is Tech job in 12 months, and estimates are obtained using Cox proportional hazards models, which include the treatment indicator and the set of controls. We estimate the model per subgroup. "Long exp. in tech" takes the value of 1 if the mentor has worked for over 5 years in the technology sector (the median value). Columns 2 and 3 describe the values for subgroups when the subgroup indicator takes the value 0, while columns 4 and 5 describe the values when the indicator takes the value 1. The last column shows the difference in the effectiveness of Mentoring between the two subgroups. Columns 1 and 3 are the average outcomes in the control groups, and columns 2 and 4 are the estimates of the average treatment effect. Standard errors are in parentheses.
Mentors with managerial experience generate 26 percentage point larger treatment effects than those without (SE 10). Similarly, mentors with over five years of technology sector experience produce 26 percentage point larger effects (SE 11). These patterns are consistent with the network channel: managers and senior professionals possess richer networks, better access to hiring decision-makers, and greater ability to facilitate job placements.
Interestingly, mentor selection patterns reveal that experienced, managerial mentors choose less promising candidates. Control group participants matched to managerial mentors have 38% baseline technology employment compared to 55% for those matched to non-managerial mentors. Yet treated participants of managerial mentors achieve higher outcomes (64% vs. 64%), because the treatment effect is substantially larger. This suggests that experienced mentors either (i.) deliberately select candidates who need more help, recognizing their own capacity to assist difficult cases, or (ii.) value different candidate attributes than baseline employment probability.
Notably, first-time mentors are as effective as repeat mentors (difference: 6pp, SE 11). This finding indicates that mentoring effectiveness stems from mentors’ accumulated professional networks, technical knowledge, and industry standing—not from mentoring skill developed through practice. A senior engineer with extensive connections generates value on her first mentoring experience; repeated mentoring does not improve outcomes.
@@
The largest heterogeneity we observe is for mothers. Among non-mothers, Mentoring generates 7 percentage point treatment effects (SE 6); among mothers, the effect is 35 percentage points (SE 10), a difference of 28 percentage points (SE 12). Mothers in the control group face substantial disadvantage, with 24% technology employment compared to 50% for non-mothers. Mentoring completely eliminates this gap: treated mothers achieve 59% employment, surpassing non-mothers.
This large effect for mothers is consistent with Mentoring addressing multiple barriers simultaneously. First, mothers may have weaker networks if career interruptions weakened professional connections. Second, work-family conflict concerns may reduce mothers’ confidence in pursuing demanding technology careers; mentor encouragement and role-modeling by successful mentors in tech may counter this.
This finding contributes to understanding gender gaps in technology beyond representation. Even among women who have acquired technical skills and are actively seeking technology employment, motherhood creates additional barriers. Traditional explanations emphasize mothers’ labor supply decisions or employer discrimination; our results suggest that confidence and network deficits also play important roles, and relatively light-touch mentoring interventions can substantially mitigate these barriers.
The contrasting heterogeneity across programs aligns with our theoretical framework. Network barriers might bind most tightly for geographically isolated women and mothers, the groups showing the largest Mentoring effects. Information frictions bind most for early-career women with limited work history, those showing the largest Challenges effects. The programs’ differential impacts across populations indicate that successful policies must target multiple, distinct barriers rather than assuming a single mechanism operates uniformly. Moreover, the patterns demonstrate that even within the selected population of women with baseline technical skills, substantial heterogeneity exists in which constraints bind most tightly, with implications for optimal program assignment that we explore in Section 6.

### v5

@@ 5.1 Treatment Effects
Two comparisons merit emphasis. First, the gap between selected-control and non-selected applicants (43% versus 19% at month thirteen) confirms that mentors identify promising candidates, whether through observable signals or information revealed in application videos. Second, the treatment effect among selected applicants (52% minus 43% = 9 percentage points) is considerably smaller than this selection effect (24 percentage points), underscoring the importance of randomization for isolating program impacts from mentor selection.
The right panel presents results for the Challenges program, comparing three groups: treated applicants (randomly assigned to participate in the Challenges program), control applicants (randomly assigned to the control group), and late applicants (who applied after randomization closed and therefore received no program exposure). The control group and late applicants exhibit nearly identical employment trajectories, with both reaching approximately 20% by month thirteen. This rate mirrors that of non-selected applicants in the Mentoring program, suggesting a common counterfactual employment rate for women seeking to enter technology absent intervention. The treatment group, by contrast, achieves 39% employment by month fourteen, nearly double the control rate.
The timing of employment gains differs notably across programs. Mentoring participants begin finding technology jobs during the program itself, consistent with mentors actively facilitating job placement through referrals or direct assistance. Challenges participants, by contrast, show employment gains concentrated after program completion, consistent with the program’s focus on building portable credentials that participants subsequently deploy in job search.
@@
Note: Share of subjects with a tech job per group across time. The month zero is the month of application to Dare IT. In the left figure, we present outcomes of three Mentoring groups: the treatment group (in blue), the control group (in green), and the applicants who were not selected by mentors (in red). In the right figure, we group focus on the Challenges program groups. The average outcome in the treated group (in blue), the control group (in green) and the late applicants to the program who were not included in the randomization and the main analysis (in red).
Table 2 presents estimated average treatment effects using two approaches: difference-in-means estimation (columns 1–3 and 5–7) and Cox proportional hazard models (columns 4 and 8). We report effects on three outcomes: any new job, new technology job, and new non-technology job. Appendix F verifies the proportional hazards assumption; Appendix C presents Kaplan-Meier survival curves confirming the timing patterns visible in Figure 2; Appendix I estimates treatment effects using causal survival forests, the method employed in the counterfactual policy analysis of Section 6; Appendix H reweighs the estimates for Challenges by stratum used for randomization.
Both programs generate substantial and statistically significant increases in technology employment. Mentoring raises the probability of technology employment by 15.3 percentage points (SE 5.7), representing a 34% increase relative to the control group mean of 45.6%. Challenges produces an 11.3 percentage point effect (SE 4.8), a 42% increase relative to the control mean of 26.6%. Cox proportional hazard estimates, which exploit the full monthly panel structure, yield nearly identical results: 14.5 percentage points for Mentoring and 11.2 percentage points for Challenges.
Table 2: Average Treatment Effects on Employment Outcomes
@@
Difference-in-means Cox PH Difference-in-means Cox PH
New job Tech job Non-tech job Tech job New job Tech job Non-tech job Tech job
@@
(0.057) (0.057) (0.031) (0.055) (0.051) (0.048) (0.042) (0.051)
ATE / Control mean (%) 13.6 33.7** -65.6** 31.7*** 14.5 42.6** -17.6 42.2**
(9.9) (14.8) (26.7) (12.0) (10.3) (18.1) (18.0) (19.3)
Treated 151 151 151 2,265 166 166 166 2,656
Control 147 147 147 2,205 218 218 218 3,488
Notes: Average treatment effects on employment outcomes measured at month 12. Columns 1–3 and 5–7 report difference-in-means estimates; columns 4 and 8 report Cox proportional hazard estimates with the full covariate set. Cox models use person-month observations. “ATE / Control mean” expresses the treatment effect as a percentage of the control group mean. Standard errors in parentheses. *** $p<0.01$, ** $p<0.05$, * $p<0.1$.
@@
Estimates in Column 3 of Table 2 suggest that Mentoring operates partly through increased job search selectivity. While Mentoring increases overall employment by only 7.8 percentage points (SE 5.7), statistically insignificant, it increases technology employment by 15.3 percentage points and decreases non-technology employment by 7.6 percentage points (SE 3.1). This implies that approximately half of Mentoring’s impact reflects participants becoming more selective, holding out for technology positions rather than accepting readily available employment in other sectors. In contrast, Challenges shows no significant effect on non-technology employment (-4.1pp, SE 4.2), consistent with this program operating through skill signaling rather than changing search behavior or persistence. The differential patterns support distinct mechanisms: Mentoring increases both participants’ capability to access technology jobs (through networks and tacit knowledge) and their willingness to persist in pursuing them (for example, through confidence-building), while Challenges primarily improves application success conditional on applying.
@@
In Appendix E, we show the results based on the outcomes survey. Response rates were modest, which is why we rely on LinkedIn data for our primary analysis. Among respondents, both treatment groups report a higher number of job offers and higher salaries than the control group. The differences between treatment and control are greater in the case of Mentoring than Challenges.
5.2 Impact of Program Participation on LinkedIn Connections
@@
We directly test the network mechanism by examining whether program participation increases professional network size, measured by LinkedIn connection counts. This analysis distinguishes network expansion from other channels: if Mentoring operates solely through skill development or signaling, we should observe no differential effect on connections relative to Challenges. Conversely, if networking is a primary mechanism, Mentoring should increase connections while Challenges, which operates entirely online without mentor introductions, should not.
@@
The data support the network mechanism. Mentoring increases LinkedIn connections by 52.7 (SE 20.8) from a control group mean of 225, a 23% increase. Challenges, by contrast, produces a statistically insignificant change of 11.4-11.4 connections (SE 17.8) from a control mean of 232. The differential effect between programs is 64 connections (SE 28), providing direct evidence that Mentoring expands professional networks while Challenges does not.
@@
Table 3 presents quantile regression estimates to examine where these effects appear in the connection distribution. Mentoring generates large, statistically significant increases at the 25th percentile (62 connections, SE 27) and median (92 connections, SE 35), but no detectable effect at the 75th percentile. The program thus particularly benefits participants with initially sparse networks. Challenges shows no significant effects at any quantile, confirming that this program does not operate through network channels.
Table 3: Treatment Effects on LinkedIn Professional Networks
Mentoring Challenges Difference
25th percentile 62** 5 57*
@@
Notes: Quantile regression estimates of treatment effects on LinkedIn connection counts. The difference column reports Mentoring minus Challenges treatment effects. Standard errors in parentheses. *** $p<0.01$, ** $p<0.05$, * $p<0.1$.
The magnitude of network expansion is economically meaningful. A 92-connection increase at the median represents substantial growth, comparable to network size differences between mid-career and senior professionals in our sample. Moreover, the measured effects likely understate true network expansion due to LinkedIn’s 500-connection display cap: 25% of Mentoring treatment participants hit this ceiling compared to 23% of controls, attenuating our estimates toward zero.
This network expansion is not driven by new employment. Both programs increase technology employment at substantial rates, yet only Mentoring participants expand their networks. If job acquisition were the primary driver of connection growth, through adding new colleagues and industry contacts post-employment, we should observe comparable network increases in both programs. The small and statistically insignificant estimate for Challenges, despite a 38% technology employment rate among participants, indicates that Mentoring’s network effects operate independently of job acquisition. This pattern is consistent with mentors actively facilitating introductions during the program, as reported in qualitative interviews, rather than networks expanding passively after job placement.
@@ 5.3 Heterogeneous Treatment Effects
The contrasting mechanisms underlying Mentoring and Challenges generate testable predictions about which participants benefit most from each program. If Mentoring operates primarily through network expansion, as suggested by the LinkedIn analysis, effects should concentrate among participants with weak existing networks. If Challenges works through skill signaling, effects should be largest for participants with weaker observable credentials. We estimate heterogeneous treatment effects using Cox proportional hazard models with the full covariate set, estimating separate models for subgroups defined by pre-treatment characteristics. [^9] The characteristics we examine are motivated by our theoretical framework. However, one might be concerned that these subgroup analyses reflect specification search. Section 6 addresses this concern using data-driven methods: we estimate optimal treatment assignment policies via causal survival forests with sample splitting, allowing the algorithm to identify predictive heterogeneity without researcher input on which variables matter. The resulting policies generate substantial gains over random assignment, confirming that policy-relevant heterogeneity exists in the data.
@@
Table 4 presents results across three participant characteristics that map to credential strength and network access. The clearest differential pattern emerges from professional experience. Challenges generates 28 percentage points larger effects for participants with under five years of experience compared to those with longer careers (SE 10), whereas Mentoring shows the opposite pattern, with 20 percentage points larger effects for experienced participants (SE 11). The cross-program differential is 49 percentage points (SE 11). This contrast directly supports our theoretical predictions: early-career women lack established work history to credibly signal technical abilities, making the Challenges portfolio particularly valuable as a complement to sparse employment records. Experienced workers, by contrast, have credentials but may lack networks in technology; mentoring addresses this constraint.
Educational background reveals an asymmetric pattern. Among Mentoring participants, those without STEM degrees show 19 percentage points larger effects than those with STEM backgrounds (SE 11). A STEM degree might bundle both network access, through connections formed during technical education, and credential signaling value; we cannot cleanly separate these channels. However, the fact that Challenges shows no educational heterogeneity (difference: 0.3pp, SE 12) is informative. One explanation is that STEM education, as measured here, is too broad to function as a meaningful signal for specific technology roles: a degree in physics or electrical engineering may not credibly signal competence as a front-end developer. If so, the Challenges portfolio provides equally valuable skill demonstration for STEM and non-STEM participants alike.
@@
Geography provides additional evidence on network mechanisms. Mentoring generates 26 percentage point effects for participants currently residing in small towns (SE 10) compared to only 3 percentage points for Warsaw residents (difference: 23pp, SE 14, $p=0.11$). Network barriers bind more tightly outside major urban centers, where local technology ecosystems are sparse. Mentor networks based in Warsaw and other major cities provide greater marginal value when participants lack local access to technology professionals. Challenges shows no significant geographic heterogeneity (difference: 2-2pp, SE 12), consistent with online portfolio signals functioning independently of applicant location.
Table 4: Heterogeneous Treatment Effects: Mentoring vs. Challenges
@@
Control mean CATE Control mean CATE Difference
@@
Yes 0.538 (0.057) 0.038 (0.079) 0.283 (0.059) 0.120 (0.099) -0.081 (0.082)
No 0.362 (0.058) 0.228*** (0.070) 0.259 (0.035) 0.116** (0.057) 0.111 (0.068)
Difference 0.176** (0.081) -0.189* (0.106) 0.024 (0.068) 0.003 (0.114) -0.193* (0.106)
Professional experience
More than 5 years 0.412 (0.049) 0.202*** (0.066) 0.310 (0.039) -0.003 (0.065) 0.205*** (0.063)
Less than 5 years 0.556 (0.075) 0.000 (0.087) 0.190 (0.042) 0.281*** (0.075) -0.281*** (0.086)
Difference -0.144 (0.089) 0.203* (0.109) 0.120** (0.057) -0.284*** (0.099) 0.486*** (0.106)
Location
Small town 0.429 (0.085) 0.260** (0.101) 0.256 (0.067) 0.149 (0.130) 0.111 (0.108)
Warsaw 0.447 (0.082) 0.033 (0.103) 0.182 (0.044) 0.166* (0.089) -0.133 (0.093)
Difference -0.019 (0.118) 0.227 (0.144) 0.074 (0.157) -0.019 (0.118) 0.246 (0.197)
Notes: Heterogeneous treatment effects on technology employment by participant characteristics. Estimates from Cox proportional hazard models with the full covariate set, estimated separately by subgroup. “Control mean” is the employment rate in the control group for each subgroup. “CATE” is the average treatment effect conditional on being in a subgroup. “Difference” column reports Mentoring CATE minus Challenges CATE. Standard errors in parentheses. *** $p<0.01$, ** $p<0.05$, * $p<0.1$.
Additional Heterogeneity in Mentoring.
The Mentoring registration survey collected more extensive information than the Challenges survey, and we observe mentor characteristics. Table 5 presents heterogeneity along these additional dimensions.
Mentor characteristics strongly predict treatment effects. Mentors with managerial experience generate 26 percentage points larger effects than those without (SE 10). Similarly, mentors with over five years of technology sector experience produce 26 percentage points larger effects than less experienced mentors (SE 11). These patterns are consistent with the network channel: managers and senior professionals possess richer networks, better access to hiring decision-makers, and greater ability to facilitate job placements directly.
Interestingly, mentor selection patterns reveal that more experienced mentors choose less promising candidates. Control group employment among participants selected by mentors with managerial experience is 38%, compared to 55% for those matched to non-managerial mentors. Yet treated participants of experienced mentors achieve substantially higher treatment effects (26pp versus 0pp). This pattern suggests that experienced mentors either deliberately select candidates who need more assistance, recognizing their capacity to help difficult cases, or value candidate attributes that differ from baseline employment probability.
First-time mentors are as effective as repeat mentors (difference: 6pp, SE 11). This finding indicates that mentoring effectiveness stems from mentors’ accumulated professional networks and industry standing rather than mentoring skill developed through practice.
Table 5: Heterogeneous Treatment Effects in Mentoring: Additional Characteristics
Yes No
Control mean CATE Control mean CATE Difference
@@
Mother 0.237 (0.070) 0.348*** (0.097) 0.532 (0.048) 0.071 (0.062) 0.277** (0.115)
Applied before 0.500 (0.087) 0.139 (0.109) 0.442 (0.047) 0.143** (0.060) -0.004 (0.124)
@@
Managerial experience 0.378 (0.056) 0.259*** (0.073) 0.549 (0.059) 0.002 (0.074) 0.257** (0.104)
Long experience in tech 0.390 (0.063) 0.277*** (0.085) 0.512 (0.054) 0.021 (0.066) 0.257** (0.107)
First-time mentor 0.444 (0.062) 0.155* (0.080) 0.476 (0.055) 0.100 (0.073) 0.056 (0.108)
Notes: Heterogeneous treatment effects in the Mentoring program by mentee and mentor characteristics. Estimates from Cox proportional hazard models with the full covariate set, estimated separately by subgroup. “Long experience in tech” indicates mentor has worked over 5 years in the technology sector (median value). Standard errors in parentheses. *** $p<0.01$, ** $p<0.05$, * $p<0.1$.
@@
The largest heterogeneity we observe is by parental status. Among non-mothers, Mentoring generates a 7 percentage point treatment effect (SE 6); among mothers, the effect is 35 percentage points (SE 10)—a difference of 28 percentage points (SE 12). Mothers in the control group face substantial disadvantage, with 24% technology employment compared to 53% for non-mothers. Mentoring eliminates this gap: 59% of treated mothers secure technology employment.
This large effect for mothers is consistent with Mentoring addressing multiple barriers simultaneously. Mothers may have weaker professional networks if career interruptions reduced opportunities to maintain connections. Work-family conflict concerns may also reduce mothers’ confidence in pursuing demanding technology careers; mentor encouragement and role-modeling by successful women in technology may counter these concerns.
This finding contributes to understanding gender gaps in technology beyond representation. Even among women who have acquired technical skills and are actively seeking technology employment, motherhood creates additional barriers. Traditional explanations emphasize mothers’ labor supply decisions or employer discrimination; our results suggest that network deficits and confidence also play important roles, and mentoring interventions can substantially mitigate these barriers.
The contrasting heterogeneity patterns across programs align with our theoretical framework. Expanding professional networks creates the greatest opportunity for geographically isolated women and mothers, the groups showing largest Mentoring effects. Building credible skill signals offers the greatest opportunity for early-career women with limited work history, those showing largest Challenges effects. These differential impacts indicate that effective policy must target multiple, distinct barriers rather than assuming a single mechanism operates uniformly. Moreover, the patterns demonstrate that even within the selected population of women with baseline technical skills, substantial heterogeneity exists in where the greatest opportunities for improvement lie. We explore implications for optimal program assignment in Section 6.